OpenStudy (anonymous):

How do I find the instantaneous slope at a point. I just need a brief explanation.

OpenStudy (wolf1728):

Take the derivative at that point.

OpenStudy (mrnood):

If the line is a straight line then the slope is constant at all points on the line. If you have an equation for the line y=f(x) then take the derivative (dy/dx or f'(x) ) Take the value of x at which you want the slope and put that value in the equation for f'(x). Evaluate that equation and that is the value of the slope at that point

OpenStudy (mrnood):

for example y=f(x)=x^2 dy/dx = f'(x) = 2x at the point x=2.5 f'(2.5)= 2(2.5)= 5